find the value of b for which the given equation is exact, and then solve it using that value of b


(xy^(2)+bx^(2)y)dx+(x+y)x^(2)dy=0


1
Expert's answer
2022-03-19T02:35:50-0400

The given equation is "(xy^{2}+bx^{2}y)dx+(x+y)x^{2}dy=0".


Since the equation is exact we must have, "\dfrac{\partial M}{\partial y} = \dfrac{\partial N}{\partial x}".


Here "M = xy^{2}+bx^{2}y; N=x^{2}(x+y)"


Then, "2xy+bx^{2} = 3x^{2}+2xy \implies b = 3"


Therefore, the given equation becomes "(xy^{2}+3x^{2}y)dx+(x+y)x^{2}dy=0", which can be written as


"\begin{aligned}
\dfrac{dy}{dx} &= \dfrac{xy^2+3x^2y}{x^2(x+y)}\\
&=\dfrac{xy(y+3x)}{x^2(x+y)}\\
&=\dfrac{y(y+3x)}{x(x+y)}\\
\end{aligned}"

This is a homogeneous equation. We use substitution to find the general solution.


Put "y=vx". Then "\dfrac{dy}{dx} = v + x\dfrac{dv}{dx}". Thus,


"\begin{aligned}
v+ x \frac{dv}{dx} &= \frac{vx(vx+3x)}{x(x+vx)}\\
&= \frac{v(v+3)}{(1+v)}\\
x \frac{dv}{dx} &= \frac{v(v+3)}{(1+v)} - v\\
&= \frac{v^2+3v-v-v^2}{(1+v)} \\
x \frac{dv}{dx}&= \frac{2v}{1+v} \\
\therefore \Big(\frac{1+v}{2v}\Big)dv &= \frac{dx}{x}\\
\frac{1}{2}\frac{dv}{v} + \frac{1}{2}dv &= \frac{dx}{x}\\
\frac{dv}{v}+dv &= 2 \frac{dx}{x}\\
\text{Integrating both sides,}\\
\log v + v &= 2\log x + \log c\\
\log v + v &= \log cx^2\\
v + \log\Big(\frac{v}{cx^2}\Big) &= 0\\
\log\Big(\frac{v}{cx^2}\Big) &= -v\\
\frac{v}{cx^2} &= e^{-v}\\
\frac{y}{cx^3} &= e^{-\frac{y}{x}}\\
ye^{\frac{y}{x}}&=cx^3
\end{aligned}"